OpenStudy (anonymous):

Solve for Sin^2 x+5sinx+6=0 I need it ASAP Please :(

OpenStudy (anonymous):

let p = sin x\[\sin^2 x + 5 \sin x + 6 = 0\]\[p^2 + 5p + 6 = 0\]solve the quadratic equation

OpenStudy (anonymous):

This does not have any solution. You would believe this after you factor \[ p^2 + 5 p + 6 \] Remember that both cos(x) and sin(x) have values in the interval [-1,1]